### v1

@@ Appendix C Ranking Algorithm
The Weighted Borda Count is a rank aggregation method that combines multiple ranked lists into a single ordering. Each post receives points based on its position within each ranking dimension (e.g., relevance, recency, popularity). Higher ranks receive more points, and these points are multiplied by user-specified weights that reflect the importance of each dimension. The final score for each post is the weighted sum of its points across all dimensions, and posts are ordered accordingly. We reproduce the pseudocode for the algorithm used in Alg. 1
@@ Appendix E Interview Protocol (1st Interview)
We include here a simplified interview protocol for the both our interviews. The interviews were semi-structured; not all questions were consistently asked to participants, and interviewers had the freedom to ask other questions to dive deeper into what participants said.
@@ E.1.1. Platforms used
    • How long have you been using the platform? How much time do you spend on the platform? Which feeds do you use?
@@ F.4. Wrap-up and Meta Reflection
(a) Feed configuration page
(b) Manage Feeds page
Algorithm 1 Weighted Borda Count for Social Media Feed Ranking
0: Ranked list of posts $R$
1:Phase 1: Generate Individual Rankings
[…]
3:for each post $p\in P$ do
4:  $(include,priority)\leftarrow\text{LLM-Evaluate}(p,U)$
5:  if $include=\text{True}$ then
6:   $R_{\text{relevance}}.\text{append}((p,priority))$
7:  end if
8:end for
[…]
11:for each post $p\in P$ do
12:  $t\leftarrow p.\text{created\_at}$
13:  $R_{\text{recency}}.\text{append}((p,t))$
14:end for
15: Sort $R_{\text{recency}}$ by $t$ descending, assign ranks 1 to $|P|$
[…]
17:for each post $p\in P$ do
19:  $R_{\text{engagement}}.\text{append}((p,e(p)))$
20:end for
21: Sort $R_{\text{engagement}}$ by $e(p)$ descending, assign ranks 1 to $|P|$
22:Phase 2: Weighted Borda Aggregation
[…]
25:for each post $p\in P_{\text{eligible}}$ do
26:  $r_{r}(p)\leftarrow\text{rank of }p\text{ in }R_{\text{relevance}}$
27:  $r_{c}(p)\leftarrow\text{rank of }p\text{ in }R_{\text{recency}}$
28:  $r_{e}(p)\leftarrow\text{rank of }p\text{ in }R_{\text{engagement}}$
30:  $\text{final\_scores}.\text{append}((p,s(p)))$
31:end for
32: Sort final_scores by score ascending
33:return $p\text{ for }(p,s)\text{ in }\text{final\_scores}]$

### v2

@@ Appendix C Ranking Algorithm
The Weighted Borda Count is a rank aggregation method that combines multiple ranked lists into a single ordering. Each post receives points based on its position within each ranking dimension (e.g., relevance, recency, popularity). Higher ranks receive more points, and these points are multiplied by user-specified weights that reflect the importance of each dimension. The final score for each post is the weighted sum of its points across all dimensions, and posts are ordered accordingly. Below, we reproduce the pseudocode for the algorithm used.
Algorithm 1 Weighted Borda Count for Social Media Feed Ranking
0: Ranked list of posts $R$
1:Phase 1: Generate Individual Rankings
[…]
3:for each post $p\in P$ do
4:  $(include,priority)\leftarrow\text{LLM-Evaluate}(p,U)$
5:  if $include=\text{True}$ then
6:   $R_{\text{relevance}}.\text{append}((p,priority))$
7:  end if
8:end for
[…]
11:for each post $p\in P$ do
12:  $t\leftarrow p.\text{created\_at}$
13:  $R_{\text{recency}}.\text{append}((p,t))$
14:end for
15: Sort $R_{\text{recency}}$ by $t$ descending, assign ranks 1 to $|P|$
[…]
17:for each post $p\in P$ do
19:  $R_{\text{engagement}}.\text{append}((p,e(p)))$
20:end for
21: Sort $R_{\text{engagement}}$ by $e(p)$ descending, assign ranks 1 to $|P|$
22:Phase 2: Weighted Borda Aggregation
[…]
25:for each post $p\in P_{\text{eligible}}$ do
26:  $r_{r}(p)\leftarrow\text{rank of }p\text{ in }R_{\text{relevance}}$
27:  $r_{c}(p)\leftarrow\text{rank of }p\text{ in }R_{\text{recency}}$
28:  $r_{e}(p)\leftarrow\text{rank of }p\text{ in }R_{\text{engagement}}$
30:  $\text{final\_scores}.\text{append}((p,s(p)))$
31:end for
32: Sort final_scores by score ascending
33:return $p\text{ for }(p,s)\text{ in }\text{final\_scores}]$
@@ Appendix E Interview Protocol (1st Interview)
We include a simplified interview protocol for the both our interviews. The interviews were semi-structured; not all questions were consistently asked to participants, and interviewers had the freedom to ask other questions.
@@ E.1.1. Platforms used
    • How long have you been/How much time do you spend on the platform? Which feeds do you use?
@@ F.4. Wrap-up and Meta Reflection
Appendix G Prompts for Scoring
We also provide the prompt used below for the curation step of the pipeline to evaluate a post’s relevance. This score, combined with the engagement and recency scores (described exactly in the algorithm), is combined through the Weighted Borda Count algorithm to obtain the final ranking.
EVALUATION GUIDELINES:
1. NEVER SHOW items should be completely blocked - if the post relates to
these topics at all,
exclude it
2. SHOW LESS OFTEN items should be filtered out unless they’re
particularly relevant or
high-quality
3. SHOW MORE OFTEN items should be given preference - include these posts
even if they’re
borderline relevant
4. ALWAYS SHOW items should be prioritized - include these posts if they
relate to the topic at all
5. NEUTRAL CONTENT (doesn’t match preferences but not excluded) should be
INCLUDED LIBERALLY for feed diversity
CONTENT MATCHING:
- Consider the main topic, theme, and context of the post
- Pay attention to hashtags, mentions, and embedded content
- Look for indirect references and related concepts
- Be more strict with exclusions (NEVER/LESS) and more lenient with
preferences (MORE/ALWAYS)
- BE INCLUSIVE of neutral content that doesn’t match
preferences but isn’t excluded
PRIORITY SCORING (for included posts):
- 10: Perfect match for ALWAYS SHOW preferences (exact topic, high
quality)
- 8-9: Good match for ALWAYS SHOW preferences (related topic, good
quality)
- 6-7: Perfect match for SHOW MORE OFTEN preferences (exact topic)
- 4-5: Good match for SHOW MORE OFTEN preferences (related topic)
- 3-4: Neutral/general interest content (no strong preference match) -
INCLUDE THESE FOR DIVERSITY
- 2-3: Borderline relevant but acceptable content -
INCLUDE THESE FOR DISCOVERY
- 1: Barely relevant but not harmful content - INCLUDE THESE FOR VARIETY
DECISION PROCESS:
1. First check if the post matches any NEVER SHOW exclusions - if yes,
exclude immediately
2. Then check if it matches SHOW LESS OFTEN exclusions - if yes, be very
strict about including it
3. Check if it matches any ALWAYS SHOW preferences - if yes, include it
with priority 8-10
4. Check if it matches SHOW MORE OFTEN preferences - if yes, include it
with priority 4-7
5. If no strong matches AND no exclusions match, DEFAULT TO INCLUDE with
priority 2-4 for content diversity
6. Only exclude if the content is clearly harmful, spam, or matches
exclusions
CONTENT DIVERSITY PRINCIPLE:
- The user wants a diverse feed that includes content beyond their
specific preferences
- If content doesn’t match exclusions, lean toward including it for
discovery and variety
- Neutral content helps prevent echo chambers and exposes users to new
topics
- Only exclude content that is clearly unwanted, not just unrelated to
preferences
RESPONSE FORMAT:
- If excluding: respond with ’no’
- If including: respond with ’yes:PRIORITY’ where PRIORITY is a
number 1-10
- Examples: ’yes:10’, ’yes:7’, ’yes:3’, ’no’
- When in doubt about neutral content, choose ’yes:3’ rather than ’no’
’yes:X’ = include this post in the user’s feed with priority X
’no’ = exclude this post from the user’s feed
Appendix H Planning Agent Details
In the planning phase, an LLM agent helps identify relevant feeds, starter packs, and lists for a user query. The agent operates in an action–observation loop: it first generates keywords from the user prompt, then queries a database via keyword matching over titles and descriptions. After proposing an initial set of sources, the user refines their preferences through the feed configuration interface. The agent then reruns the same keyword-generation and retrieval process using the updated preferences, and the user reviews and selectively includes additional suggested sources before proceeding to feed generation. We provide the prompt provided to the agent to generate keywords below:
Based on these preferences, generate 5-8 SHORT SINGLE-WORD search
terms that would help find relevant feeds, starter packs, and
lists in a content discovery database.
CRITICAL: Always start with the most OBVIOUS, BASIC terms from
their preferences:
- If they mention "dogs and cats" → include "dogs", "cats", "pets"
- If they mention "machine learning" → include "machine",
"learning", "ai", "ml"
- If they mention "Love Island US" → include "love", "island", "reality"
- If they mention "cooking" → include "cooking","recipes", "food"
Then add 2-3 related terms for broader discovery.
Return a JSON object with:
- "search_terms": array of 5-8 single words starting with the
most obvious terms
- "reasoning": brief explanation of why these terms were chosen
Example for "I want posts about cute dogs and cats":
"search_terms": ["dogs", "cats", "pets", "animals", "cute",
"adorable"],
"reasoning": "Start with obvious terms dogs/cats/pets, then
add related animal and emotion terms"
IMPORTANT:
- ALWAYS include the most obvious, basic words from their preferences
FIRST- Don’t overcomplicate - use simple, direct terms
- Use only single words for better database matching
Appendix I Generative AI Disclosure